OpenStudy (anonymous):

y+2x=5; solve for x=-1,0,3

OpenStudy (anonymous):

Help anyone??????????????

OpenStudy (anonymous):

i'm not sure what you mean, i think that the second part means that x has multiple values, so based on the information that you have given, i imagine that you are solving for y since there is no information for y given. y+2(-1)=5 y-2=5 (y1 = 7) y+2(0)=5 (y2 = 5) y+2(3)=5 y+6=5 (y3 = -1) ***I hope that that is what you are looking for!! Good luck!!!*** Medal????

OpenStudy (anonymous):

For this problem you have to make the equation into y=mx+b format, then plug in the x- values.

OpenStudy (anonymous):

xould i get a medal please???